OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)?

OpenStudy (ivettef365):

first you need to find the slope, do you know how to

OpenStudy (anonymous):

yes y1 + y2 then what

OpenStudy (anonymous):

@amistre64

OpenStudy (texaschic101):

m = y2 - y1 / x2 - x1) is the formula to find the slope. set 1 (-2,3)....set 2 (2,7) lets sub in... m = (7 - 3)/(2 - (-2) m = 4/4 = 1 now use the point slope formula : y - y1 = m(x - x1) I am using points (2,7)....you could also use (-2,3)...either one will work slope (m) = 1 y - y1 = m(x - x1) y - 7 = 1(x - 2) y - 7 = x - 2 y = x -2 +7 y = x + 5 <-- this is your equation